Solve for x: 4x/x-3= 2 + 12/x-3
Answer by mananth(16946)

4x/x-3= 2 + 12/x-3
4x/x-3= 2*(x-3)+12 / x-3
4x(x-3)/(x-3) =(2x-6 +12)
4x= 2x-6+12
4x-2x=6
2x=6
x=3